OpenStudy (precal):

slopefields general question if I am given a dy/dx equation how do I match it up to the slopefield

OpenStudy (anonymous):

Plug in points \((x,y)\) of your choice. For example, if \(\dfrac{dy}{dx}=xy\), you can plug in \((0,0)\) and you have \(\dfrac{dy}{dx}=0\). If you pick \((1,1)\), \(\dfrac{dy}{dx}=1\). And so on.

OpenStudy (mathteacher1729):

Look for lines of equal slope on the slopefield. "Where is dy/dx = 0 ? Where is dy/dx > 0?, Where is dy/dx < 0? Where is dy/dx not defined? etc.

OpenStudy (precal):

Thanks not sure this is what I was looking for. But I am working some problems out using slopefields.

OpenStudy (mathteacher1729):

What are the problems asking you to do?

OpenStudy (precal):

They are misc. AP problems. Just stuck on them.

OpenStudy (precal):

They are just really awful problems. I finally figured out how to use my TI-89